OpenStudy (anonymous):

Which mixed number is greater than 6.78?

OpenStudy (tester97):

Express 6.78 as a fraction. Obviously it's going to be 6 and something, so what's .78 as a fraction? There are 2 digits to the right of the decimal point, so that means 0.78 it's out of hundreths. 0.78 = 78/100, which simplifies to 39/50. So 6.78 = 6 39/50 6 2/3 = 6.6666, so obviously this is less. The answer is not A. 6 3/4 = 6.75, so again, this is less. The answer is not B. 6 7/9 = 61/9 = 6.7777777777777777 so again this is just less. The answer is not C. The answer must therefore be D, but to prove this: 6 4/5 = 34/5 = 6.8 So yes, the answer is D.